OpenStudy (sasogeek):

evaluate \[\huge \sqrt{ \frac{0.81 \times 10^{-5}}{2.25 \times 10^{7}}} \]

OpenStudy (anonymous):

sorry i can't see

OpenStudy (jamesj):

Hint: \( 0.9^2 = 0.81 \) and \( 1.5^2 = 2.25 \).

OpenStudy (sasogeek):

tomas

OpenStudy (bahrom7893):

and 10^(-5)/(10^7) = 10^(-12)

OpenStudy (anonymous):

Divide 0.81 by 2.25 so you get 0.36 0.36x10^-7x10^-5 0.36x10^-12 take full square root using calculator: you get 6x10^-7 which is the answer.

OpenStudy (sasogeek):

well i'm not supposed to use a calculator so how do i use my head to arrive at 0.36... it seems possible but not easy -_-

OpenStudy (jamesj):

Use the hint I wrote above. Then your expression is equal to \[ \sqrt{ \frac{ 0.9^2}{1.5^2} \times 10^{-12} } \] \[ = \frac{9}{15} \times 10^{-6} \] \[ = \frac{3}{5} \times 10^{-6} \] \[ = 0.6 \times 10^{-6} \] \[ = 6 \times 10^{-7} \]

OpenStudy (anonymous):

@Saso-Take 0.81 as 81 and 2.25 as 225 then use long division method.

OpenStudy (anonymous):

the hard part is probably seeing that \[\sqrt{.81}=.9\] this is because \[.9^2=.81\] but also because \[.91=\frac{81}{100}\] so \[\sqrt{.81}=\sqrt{\frac{81}{100}}=\frac{9}{10}=.9\]